Calculate volume of H2 gas kept at STP if it contains as many H atoms as in 98 gm H3PO4

Grade:

1 Answers

Arun
5 years ago
Molar mass of the Phosphoric acid = 98 g/mole. 
Mass of the Phosphoric acid = 98 g. 
∴ No. of molecules of the Phosphoric acid = 6.022 × 10²³ molecules.
Now, 
1 molecule of the Phosphoric acid have 3 atoms of Hydrogen. 
∴ 6.022 × 10²³ molecules of Phosphoric acid have 6.022 × 10²³ × 3 atoms of Hydrogen = 1.81 × 10²⁴ atoms. 
Now, Acc. to the question, 
Hydrogen gas at S.T.P. also have the same number of atoms, i.e., 1.81 × 10²⁴ atoms. 
∴ Mass of the Hydrogen containing 1.81 × 10²⁴ atoms. = Molar Mass × 1.81 × 10²⁴/Avogadro's number. 
= 2 × 1.81 × 10²⁴/6.022 × 10²³
= 2 × 3
= 6 g. 

We know that the 1 mole of any gas at S.T.P. occupies the volume of the 22.4 liter. 
∴ 2 g of the Hydrogen at S.T.P. occupies the volume of the 22.4 liter. 
∴ 6 g of the Hydrogen at S.T.P. occupies the volume of the 22.4 × 6/2 liter. 
 = 22.4 × 3 = 67.2 liter. 

Hence, the volume of the Hydrogen at S.T.P. is 67.2 liter.